OpenStudy (anonymous):

To solve the system of equations by elimination, x + 3y = 3 and -2x + 4y = 6, what number do you multiply the first equation by to eliminate the x? A. 1 B. 0 C. 3 D. 2

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

@AntoinetteWillman Are you still there?

OpenStudy (anonymous):

I was thinking 3.

OpenStudy (anonymous):

D. 2 makes it 2x so you can cancel out 2(x+2y)=3(2) 2X+4y = 6 2X+4y = 6 +(-2x+4y=6) ______________ 8y=12 Solve.

OpenStudy (anonymous):

A?